OpenStudy (anonymous):

which of the following are equal to the expression below ? check all that apply 2* sqrt 25 a. 50 b .100 c.25 d.sqrt 100 e.10 f. sqrt 50 ?

OpenStudy (anonymous):

I'm sure that if you'd only TRY to solve it... you would do it with in 10 seconds... but you didn't even try.. I'm sure !!

OpenStudy (anonymous):

2 times 25 is 5o correct ?

OpenStudy (anonymous):

\[\LARGE \sqrt{25}=?\]

OpenStudy (anonymous):

50

OpenStudy (anonymous):

whoaa... that's one UGLY answer ! why do you think 50 ??

OpenStudy (anonymous):

10 or 100 ?

OpenStudy (anonymous):

\[\LARGE \sqrt{25}=\sqrt{5^2}=??\]

OpenStudy (anonymous):

none of them ! ...

OpenStudy (anonymous):

\[\LARGE \sqrt{25}=\sqrt{5^2}=\left(5\right)^{\frac22}=??\]

OpenStudy (anonymous):

never mind... the answer is e